OpenStudy (anonymous):

PLEASE!!!! I NEED HELP! ASAP!!!!! I DON'T GET IT!! Determine whether the given ordered pair (1/4, 3/4) is a solution of the system. 5x+3y=-1 x-y=1

OpenStudy (anonymous):

apply your numbers to the equations and see if they are true.

OpenStudy (turingtest):

put the numbers where x and y go in the equations and see if it's true\[5(1/4)+3(3/4)=^?1\]etc.

OpenStudy (anonymous):

okay and then wat? do i do distributive property??

OpenStudy (turingtest):

\[5(1/4)+3(3/4)=5/4+9/4=14/4\neq-1\]so not a solution.

OpenStudy (anonymous):

oh okay! i get it now! Thanks so much!!